# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ Abstract
This paper establishes a fundamental impossibility theorem: no LLM capable of performing non-trivial knowledge aggregation can simultaneously achieve truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality. The impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
We establish this result by describing the inference process as an auction of ideas, where distributed components compete exploiting their partial knowledge to shape responses. The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity). In particular, we show how to quantify the creation of overconfident or intuitive responses–the signature of both hallucination and creativity, or imagination.
To support this analysis, we introduce the complementary concepts of the semantic information measure and the emergence operator to model bounded reasoning in a general setting. We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, the idealized unconstrained reasoning strictly preserves semantic content.
By demonstrating that hallucination and imagination are mathematically identical phenomena–grounded in departures from truthfulness, semantic information conservation, revelation of relevant knowledge, and knowledge-constrained optimality–we offer a principled foundation for managing these behaviors in advanced AI systems. Finally, we present some speculative ideas to inspire evaluation and refinements of the proposed theory.
@@ 1 Introduction
Large language models have been famous for producing content that is factually incorrect, inconsistent, or well fabricated and compelling–a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, like mixture-of-experts (MoE), retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
@@
The idea of hallucination has clearly much to do with creativity. Therefore, whether it is desired or not depends on circumstances. We could argue that a much better name for that phenomenon we are dealing with may be imagination. As an essential component of intelligence, imagination is responsible for seeing into the future and learning from the past. It is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened. It is one of the driving forces of art and science, and exploration of the unknown. But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will all the compelling visions of reality that does not exist.
@@
This paper shows that hallucination is a manifestation of fundamental limitations of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality. Hallucinations will manifest one way or the other.
@@
In this paper we show the impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraints. We model inference as an auction of ideas where components encoding knowledge–attention heads, circuits, and activation patterns in LLMs–compete to influence response generation by bidding with their information contributions.
@@
This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory to establish impossibility first in an idealized setting where knowledge is independently distributed. We then extend beyond that idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules–the standard training paradigm for modern LLMs. Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful reporting inevitably creates excess confidence. Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable (Jensen) gap that quantifies the violation of the semantic information conservation principle.
@@
The progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our engineering choices but a fundamental property of information aggregation itself.
1.1 An Illustrative Example
@@ 1.2 The Main Result
The example above illustrates the core insight of this paper that we prove formally in complementary and general settings. Any response to a non-trivial query sacrifices some aspects of an ideal hallucination-free response.
First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground-truth matching. Second, we go beyond that idealized model and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers. This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
@@
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge. And so on and so forth, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
In other words and on the other hand, creative tasks necessarily involve hallucination, or better yet, imagination. When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality) or generate plausible imaginary extensions (violating conservation). Creativity requires violation of information conservation. Should we control such creative insights boldly going where no one has gone before?
Despite its name, the impossibility theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights. The theorem suggests that hallucination is an inherent feature of any capable inference systems. Since it is inevitable, it should be managed. Different applications demand different trade-offs: medical diagnosis systems might maximize truthfulness at the cost of creativity, while creative writing tools might embrace controlled hallucination for narrative richness. Understanding these trade-offs enables principled design.
@@ 1.3 Paper Organization
The remainder of this paper is organized as follows. Section 2 reviews the existing literature on LLM hallucinations, situating our contribution within theoretical, mechanistic, and empirical contexts. Section 3 summarizes our main contributions.
Section 4 introduces the rigorous mathematical foundations of our analysis. We define the semantic information measure ($\mu_{C}$) and the emergence operator ($\mathcal{E}_{C}$) to model bounded reasoning, establish the principle of information conservation under unlimited reasoning, and formalize the auction-theoretic model of inference. We also define the four essential properties required for hallucination-free mechanisms.
Section 5 presents the central Impossibility Theorem, offering the first two proofs. Proof I (Theorem 6) utilizes mechanism design theory in an idealized setting of independent private knowledge. Proof II (Theorem 8) generalizes the result to probabilistic settings using the theory of proper scoring rules.
Section 6 connects these theoretical results to concrete LLM implementations. We analyze the transformer architecture and present Proof III (Theorem 9), demonstrating how the Log-Sum-Exp structure inherently violates information conservation and quantifying the resulting Jensen gap ($\Gamma$).
Finally, Section 7 discusses the implications of the theorem, exploring strategies for managing trade-offs, unifying hallucination and imagination, and speculating on the connections to fundamental limits of reason and models of consciousness. Section 8 concludes the paper.
@@ 2 Related Work
The conclusion that LLMs hallucinate is not new but the evaluation of their scope and underlying justifications differ significantly from what we offer here. To our best knowledge, the impossibility theorem presented in this paper offers a novel first-principles explanation for the phenomenon, distinct from existing approaches. This section situates our contribution by reviewing three key areas: theoretical arguments for the inevitability of hallucination, mechanistic studies into their origins, probabilistic frameworks for their detection, and empirical mitigation techniques.
@@ 2.1 Theoretical Perspectives on Inevitability
In their work, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from their fundamental mathematical and logical structure. Their conclusion is that LLMs cannot detect and eliminate hallucinations by themselves and therefore must hallucinate. The analysis draws on Goedel’s incompleteness theorem and Turing’s undecidability theorem. First, incompleteness argument is called to show that an LLM cannot verify its responses because of the training set incompleteness. Second, the hallucination (self)detection problem is reduced to the halting problem, which is undecidable. While the finale conclusion aligns with that of our paper, the arguments presented are more about the nature of hallucination detection with limited ground-truth access (see Section 4.2 for the ground-truth concept discussion) than the inference process limitations. In other words, it is about practical impossibility of real-time, exhaustive fact-checking.
@@ 2.3 Probabilistic Frameworks
Together, these complementary approaches establish hallucination probability arising from multiple independent mathematical constraints.
@@ 2.4 Empirical Mitigation Techniques
Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails. Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination–an observation our theorem explains from first principles. Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
@@ 3 Main Contributions
We introduce three major contributions that expand our understanding of (neural) inference.
@@
The Auction of Ideas: We formalize LLM inference as a mechanism design problem where neural components (attention heads, circuits, activation patterns) are agents competing with private knowledge to influence response generation. This novel perspective enables us to apply powerful results from game theory. To our best knowledge, that is one of the first applications of mechanism design framework in that setting. Also, that framework relates surprisingly well to the multiple drafts model of consciousness proposed by Dennett in Dennett1991-DENCE and provides a mathematical justification for it.
@@
Semantic Information Theory with Computational Bounds: We introduce a rigorous mathematical framework based on two complementary concepts. The semantic information measure $\mu_{C}$, parameterized by computational budget $C$, quantifies how knowledge reduces uncertainty within bounded reasoning. The emergence operator $\mathcal{E}_{C}$ formalizes how reasoning makes latent knowledge explicit without creating it ex nihilo. In that framework we establish the principle of information conservation: unlimited reasoning reveals but never creates information. That resolves the apparent paradox of how LLMs can generate novel insights while being deterministic systems. Those insights were always latent, requiring only computational work to access.
@@
The Impossibility Proofs: We prove the same impossibility through three complementary lenses. In Theorem 6 we use mechanism design with independent private knowledge setting to establish an idealized border case. In Theorem 8 we extend the result to the general setting of correlated beliefs via proper scoring rules. Then, in Theorem 9 we demonstrate the impossibility in actual transformer implementations through the log-sum-exp Jensen gap.
@@
This multi-faceted approach proves the impossibility is not an artifact of our theoretical choices but emerges from fundamental limits of reason and language. We contribute to that bold claim in the following way.
@@
Unifying hallucination and creativity (imagination): We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
@@
Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores absolute boundaries of what intelligence can achieve. This places LLM limitations within the grand tradition of fundamental impossibility results.
@@
Experimental Philosophy Platform: The framework transforms LLMs into accessible (philosophical) laboratories for testing hypothesis of mind. By providing precise mathematical definitions of concepts like knowledge accessibility,semantic conservation, and cognitive hunger (information contribution), we enable rigorous experimental investigation of many claims.
@@
Theoretical Foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior. The semantic information measure quantifies knowledge accessibility within computational bounds. The emergence operator captures how reasoning transforms latent knowledge into accessible form. The Jensen gap provides a measurable signature of hallucination in actual systems. This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself. Previous work identified important statistical correlations or architectural features associated with hallucination. We show why these patterns must exist and how to navigate possible trade-offs.

### v7

@@ Abstract
This paper establishes a fundamental Impossibility Theorem: no LLM capable of performing non-trivial knowledge aggregations can simultaneously achieve truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality. The impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
We establish this result by describing the inference process as an auction of ideas, where distributed information processing components compete to be included in a response while exploiting accessible and relevant knowledge. The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity).
To model the inference process in a general setting, we introduce two complementary concepts, the semantic information measure and the emergence operator. These concepts then allows us to investigate the limits of computationally bounded and unbounded reasoning. While bounded reasoning can make information accessible, thereby providing valuable insights and inspirations, the unbounded reasoning strictly preserves semantic content already hidden in any set of encoded knowledge. Furthermore, we show that meaningful reasoning must violate the information conservation principle, which is a manifestation of the conservation-reasoning dichotomy.
By demonstrating that hallucination and imagination are mathematically identical phenomena—grounded in departures from truthfulness, semantic information conservation, revelation of relevant knowledge, and knowledge-constrained optimality—we offer a principled foundation for managing these behaviors in advanced AI systems. Finally, we present some speculative philosophical ideas to inspire evaluation and refinements of the proposed theory.
@@ 1 Introduction
Large language models have been famous for producing content that is factually incorrect, inconsistent, fabricated and yet compelling and immensely impressive—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, like mixture-of-experts (MoE), retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
@@
The idea of hallucination has much to do with creativity and whether it is desired or not depends on circumstances. We could argue that a much better name for the phenomenon we are dealing with may be imagination. As an essential component of intelligence, imagination is responsible for seeing into the future and learning from the past. It is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened. It is one of the driving forces of art and science, and exploration of the unknown. But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will all the compelling visions that do not exist in any training set defining the ground truth for learning.
@@
This paper shows that hallucination is a manifestation of fundamental limitation of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible to simultaneously achieve: (1) truthful response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality. Hallucinations manifest one way or the other. These properties are not arbitrary, but come from the exhaustive and rigorous game-theoretic studies of information aggregation in multi-agent systems.
@@
The impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraints. In each case, the inference process is described as an auction of ideas in which components encoding knowledge—like attention heads, circuits, and activation patterns in transformer-based LLMs—compete to influence response generation by bidding with their information contributions.
@@
That game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory to establish impossibility first in an idealized setting where knowledge is independently distributed. We then go beyond that idealized setting to encompass systems where components output probability distributions optimized through strictly convex loss functions–the standard training paradigm for modern LLMs. Here, the impossibility manifests largely (but not only) through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful reporting inevitably creates excess confidence. Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable (Jensen) gap that quantifies the violation of the semantic information conservation principle.
@@
The progression from abstract and idealized models of game theory, through probabilistic frameworks, to concrete neural architectures, demonstrates that the impossibility is not an artifact of our engineering choices but a fundamental property of information aggregation itself.
Terminology Note
This paper uses the term hallucination in its established technical sense within artificial intelligence research, referring to AI systems generating outputs that are factually incorrect, inconsistent, or unsupported by their training data. We acknowledge that this technical usage differs from the clinical term describing perceptual experiences in certain medical and psychiatric conditions. We recognize and respect the serious challenges faced by individuals experiencing clinical hallucinations and their families. Our use of this terminology follows established conventions in machine learning literature and is not intended to trivialize or misrepresent these medical experiences.
1.1 An Example
@@ 1.2 The Main Result
The example above illustrates the core insight of this paper. Any response to a non-trivial query sacrifices some aspects of an ideal hallucination-free response. Hallucinations (or creative responses) manifest as violations of at least one of those aspects.
@@
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained model will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge. And so on and so forth, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
In other words and on the other hand, creative tasks necessarily involve hallucination, or better yet, imagination. When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality of user-experience maximizing response) or generate plausible imaginary extensions (violating conservation of information by generating new ideas in the internal representation space).
There is an important distinction we must emphasize. We define hallucination as a structural violation of idealized inference principles, leading to fabricated confidence of any generative response. That is what the Impossibility Theorem proves is inevitable. What the Impossibility Theorem does not prove, is that output must be incorrect relative to ground truth. That may happen as a potential consequence of hallucination, but it does not have to happen. Excess confidence of inevitable hallucination may as well result in factually correct lucky hallucination. In other words, any truthful response revealing relevant knowledge to maximize user’s satisfaction is provably non-conservative (in information creation) regardless of its veracity.
Despite its name, the Impossibility Theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights. The theorem suggests that hallucination is an inherent feature of any capable inference system. Since it is inevitable, it should be managed. Different applications demand different trade-offs: medical diagnosis systems might maximize truthfulness at the cost of creativity, while creative writing tools might embrace controlled hallucination for narrative richness. Understanding these trade-offs enables principled design.
@@ 1.3 Paper Organization
The paper starts with theoretical foundations and progresses through mathematical proofs and examples to philosophical implications.
Section 2 situates our contribution within the broader landscape of hallucination research, reviewing theoretical perspectives on inevitability, mechanistic interpretability studies, probabilistic frameworks, and empirical mitigation techniques.
Section 3 synthesizes the key contributions: the auction-theoretic model of inference, the semantic information framework with computational bounds, the triple-proof strategy establishing impossibility across different mathematical domains, and the unification of hallucination with creativity.
Section 4 constructs the mathematical framework. It formalizes knowledge as a Polish metric space, introduces the semantic information measure $\mu_{C}$ parameterized by computational budget, and defines the emergence operator $\mathcal{E}_{C}$ that captures how reasoning makes latent knowledge explicit. It establishes fundamental results about information preservation under unlimited reasoning and defines the auction-theoretic model where neural components compete as agents with private knowledge. The section concludes by defining the four essential properties of hallucination-free mechanisms.
Section 5 presents the core impossibility results through complementary proofs. Theorem 6 applies mechanism design theory in the idealized setting of independent private knowledge. Theorem 8 extends the result to probabilistic predictions using proper scoring rules.
Section 6 bridges theory to implementation, analyzing a concrete micro-transformer example. We trace inference from embeddings through attention mechanisms to final predictions, demonstrating precisely where the impossibility emerges. Theorem 9 proves that the log-sum-exp structure promotes hallucination as excess confidence.
Section 7 examines safety implications through the lens of bounded creativity. Theorem 11 establishes that meaningful reasoning necessarily violates strict information conservation. Then it shows that while creativity must remain bounded for safety, enforcing strict conservation of information eliminates meaningful reasoning entirely. The section establishes that safety is achieved through alignment within authorized knowledge boundaries rather than through conservation alone.
Section 8 explores broader implications. We analyze escape strategies through hybrid architectures, connect our results to fundamental limits of reason alongside Gödel and Heisenberg, and provide mathematical grounding for Dennett’s Multiple Drafts model of consciousness. The section deliberately embraces speculation to inspire future research directions.
Section 9 synthesizes the journey from mathematical necessity to philosophical implications, outlining how our framework opens new avenues for understanding intelligence itself.
@@ 2 Related Work
The conclusion that LLMs hallucinate is not new but the evaluation of their scope and underlying justifications differ significantly from what we propose in this paper. To our best knowledge, the Impossibility Theorem presented here offers a novel first-principles explanation for the phenomenon, distinct from existing approaches. This section situates our contribution by reviewing three key areas: theoretical arguments for the inevitability of hallucination, mechanistic studies into their origins, probabilistic frameworks for their detection, and empirical mitigation techniques.
@@ 2.1 Theoretical Perspectives on Inevitability
In their work, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from their fundamental mathematical and logical structure. Their conclusion is that LLMs cannot detect and eliminate hallucinations by themselves and therefore must hallucinate. The analysis draws on Goedel’s incompleteness theorem and Turing’s undecidability theorem. First, incompleteness argument is called to show that an LLM cannot verify its responses because of the training set incompleteness. Second, the hallucination (self)detection problem is reduced to the halting problem, which is undecidable.
While the final conclusion aligns with that of our paper, the arguments there are more about the nature of hallucination detection with limited ground-truth access (see Section 4.2 for the ground-truth concept discussion) than the inference process limitations. In other words, it is about practical impossibility of real-time, exhaustive fact-checking.
@@ 2.3 Probabilistic Frameworks
In more recent kalai2025why , Kalai et al. expand the theory of hallucination by connecting it with the computational learning theory. They demonstrate that any language model can be viewed as an IIV-type (Is-It-Valid) classifier trained on a set of responses labeled either as valid or error. Then, it follows that the statistical objectives of pretraining necessarily induce generative errors when the underlying data distribution is complex, characterized by poor model fit, or undersampled. This establishes that hallucination arises from the statistical impossibility of perfect learning.
In contrast, we demonstrate the structural impossibility of perfect inference. For non-trivial queries or contexts, that impossibility of perfect response aggregation persists even when the statistical impossibility of perfect learning is eliminated in kalai2025why . Suppose in the training set of examples there plausible errors exist, i.e., $\min_{c}|\mathcal{E}_{c}|>0$ ensuring the classification problem is well defined and non-trivial. Suppose the statistical learning problem is solved for perfect data and model capacity, so that the positive error rate of the trained LLM is not forced by the identified fundamental bounds. Then, even if the model possesses perfect statistical knowledge, this knowledge may still be distributed across the internal states (e.g., different attention heads encoding different facets of the truth). Since the context is non-trivial and the knowledge is distributed, the beliefs of the agents bidding with ideas will necessarily diverge, and the aggregation mechanism must synthesize these divergent beliefs. This process necessarily generates excess confidence or violates the relevant knowledge revelation. As a result, the impossibility of perfect response aggregation persists even when the statistical impossibility of perfect learning is eliminated for non-trivial contexts.
Together, these complementary approaches establish that hallucination is inevitable due to complementary mathematical constraints: the limits of learning the true distribution from data and the fundamental structural limits of aggregating distributed knowledge during inference.
@@ 2.4 Empirical Mitigation Techniques
Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails. Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination–an observation our theorem explains from first principles. Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets. Finally, Dubanowska et al. dubanowska2025representation show that hallucination detection methods based on internal representations suffer from spurious correlations, and so the entire approach needs rethinking.
@@ 3 Main Contributions
This paper aims at expanding our understanding of (neural) inference processes and their limitations. That is achieved by introducing the following contributions.
@@
The Auction of Ideas: We formalize LLM inference as a mechanism design problem where neural components (attention heads, circuits, or activation patterns in transformers) are agents competing with private knowledge to influence response generation. This novel perspective enables us to apply powerful results from game theory. To our best knowledge, that is one of the first applications of mechanism design framework in that setting. Also, that framework relates surprisingly well to the multiple drafts model of consciousness proposed by Dennett in Dennett1991-DENCE and provides a mathematical justification for it.
@@
Semantic Information Theory: We introduce a rigorous mathematical framework based on two complementary concepts. The semantic information measure $\mu_{C}$, parameterized by computational budget $C$, quantifies how knowledge reduces uncertainty within computationally bounded reasoning. The emergence operator $\mathcal{E}_{C}$ formalizes how reasoning makes latent knowledge explicit without creating it ex nihilo. In that framework we establish the principle of information conservation: unlimited reasoning reveals but never creates information. That also establishes a dichotomy of reasoning and semantic information conservation.
@@
The Impossibility Proofs: We prove the same impossibility through three complementary lenses. In Theorem 6 we use mechanism design with independent private knowledge setting to establish an idealized border case. In Theorem 8 we extend the result to the general setting of correlated beliefs via proper scoring rules. Then, in Theorem 9 we demonstrate the impossibility in actual transformer implementations through the log-sum-exp Jensen gap. Finally, in Theorem 11 we show that semantic information conservation and meaningful reasoning are mutually exclusive.
@@
This multi-faceted approach proves the impossibility is not an artifact of our theoretical choices but emerges from fundamental limits of reason and language. We contribute to that bold claim in the following ways.
@@
Unifying hallucination and imagination: We argue these are mathematically identical phenomena that arise from the aggregation process that creates confidence exceeding constituent evidence.
@@
Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores limits of what intelligence can achieve. This places LLM limitations within the grand tradition of fundamental impossibility results.
@@
Experimental philosophy: The framework transforms LLMs into accessible (philosophical) laboratories for testing hypothesis of mind. By providing precise mathematical definitions of concepts like knowledge accessibility,semantic conservation, and cognitive hunger (information contribution), we enable rigorous experimental investigation of many claims.
@@
Theoretical foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior. The semantic information measure quantifies knowledge accessibility within computational bounds. The emergence operator captures how reasoning transforms latent knowledge into accessible form creating ground for new insights. This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself. Previous work identified important statistical correlations or architectural features associated with hallucination. We show why these patterns must exist and how to navigate possible trade-offs.
We should point out that empirical validation of the formal results is beyond the scope of this paper. However, we provide toy examples that can be expanded in future research on hallucination detection and artificial creativity. This empirical work is crucial for translating the theoretical insights into practical tools. See Dubanowska et al. dubanowska2025representation and Brzozowski brzozowski2025bimodality for some initial results.